# Provable Privacy with Non-Private Pre-Processing

### v1

@@ Abstract
When analyzing Differentially Private (DP) machine learning pipelines, the potential privacy cost of data-dependent pre-processing is frequently overlooked in privacy accounting. In this work, we propose a general framework to evaluate the additional privacy cost incurred by non-private data-dependent pre-processing algorithms. Our framework establishes upper bounds on the overall privacy guarantees by utilising two new technical notions: a variant of DP termed Smooth DP and the bounded sensitivity of the pre-processing algorithms. In addition to the generic framework, we provide explicit overall privacy guarantees for multiple data-dependent pre-processing algorithms, such as data imputation, quantization, deduplication and PCA, when used in combination with several DP algorithms. Notably, this framework is also simple to implement, allowing direct integration into existing DP pipelines.
@@ 1 Introduction
With the growing emphasis on user data privacy, Differential Privacy (DP) has become the preferred solution for safeguarding training data in Machine Learning (ML) and data analysis (Dwork et al., 2006). DP algorithms are designed to ensure that individual inputs minimally affect the algorithm’s output, thus preserving privacy. This approach is now widely adopted by various organizations for conducting analyses while maintaining user privacy (AppleDP, 2017; U.S. Census Bureau, 2020).
Pre-processing data is a standard practice in data analysis and machine learning. Techniques such as data imputation for handling missing values (Anil Jadhav and Ramanathan, 2019), deduplication for reducing memorization and eliminating bias (Kandpal et al., 2022; Lee et al., 2022), and dimensionality reduction for denoising or visualization (Abadi et al., 2016; Zhou et al., 2021; Pinto et al., 2024) are commonly used. These pre-processings are also prevalent prior to applying DP algorithms, a pipeline that we term pre-processed DP pipeline. Among other uses, this has been shown to improve privacy-accuracy trade-off (Tramer and Boneh, 2021; Ganesh et al., 2023).
@@
There are multiple strategies to address this. A straightforward method to derive privacy guarantees for this pipeline is to use group privacy where the size of the group can be as large as the size of the dataset, thereby resulting in weak privacy guarantees. This idea, albeit not in the context of pre-processing, was previous explored under the name Dependent Differential Privacy (DDP) (Zhao et al., 2017; Liu et al., 2016). Another approach is to use public data for the pre-processing. Broadly referred to as semi-private learning algorithms, examples of these methods include pre-training on public data and learning projection functions using the public data e.g. Pinto et al. (2024); Li et al. (2022a, b); Yu et al. (2021). Despite their success, these methods crucially rely on the availability of high-quality public data.
@@
    We introduce a generic framework to quantify the privacy loss in the pre-processed DP pipeline. Applying this framework, we evaluate the impact of commonly used pre-processing techniques such as deduplication, quantization, data imputation, and PCA on the overall privacy guarantees.
@@
Another line of research, focusing on privacy in correlated datasets (Liu et al., 2016; Zhao et al., 2017; Humphries et al., 2023), show that correlations in the datasets can increase privacy risk of ML models. In response, Pufferfish Differential Privacy (Kifer and Machanavajjhala, 2014; Song et al., 2017) and Dependent Differential Privacy Liu et al. (2016); Zhao et al. (2017) was proposed as privacy notions tailored for datasets with inherent dependencies. However, these definitions usually require complete knowledge of the datasets’ dependency structure or the data generating process and sometimes leads to vacuous privacy guarantees. Moreover, the application of these privacy notions usually complicates the privacy analysis, as many privacy axioms, such as composition, does not hold under these more general notions.
@@ 2.1 Rényi Differential Privacy
Differential privacy restricts the change in the output distribution of a randomized algorithm by altering a single element in the dataset. Formally, for $\varepsilon,\delta>0$, a randomized algorithm ${\mathcal{A}}$ satisfies $(\varepsilon,\delta)$-DP if for any two dataset $S,S^{\prime}$ that differ by exactly one element and any possible output of the algorithm $O$,
${\mathbb{P}}\left[{{\mathcal{A}}(S)\in O}\right]\leq e^{\varepsilon}{\mathbb{P}}\left[{{\mathcal{A}}(S^{\prime})\in O}\right]+\delta.$ (1)
In this work, we mainly focus on a stronger notion of DP, known as Rényi Differential Privacy (RDP)Mironov (2017), based on Rényi divergence between two distributions.
@@
Let $P,Q$ be two probability distribution with $supp(P)\subseteq supp(Q)$. Let $\alpha>1$. The Rényi divergence with order $\alpha$ between $P$ and $Q$ is defined as
$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log{\mathbb{E}}_{x\sim Q}\left({\frac{P(x)}{Q(x)}}\right)^{\alpha}.$
@@
Let $\varepsilon(\alpha)$ be a function that maps each $\alpha$ to a positive real number. A randomized algorithm ${\mathcal{A}}$ is $(\alpha,\varepsilon(\alpha))$-RDP if for all $\alpha>1$, for any neighboring datasets $S,S^{\prime}$ with $d_{H}(S,S^{\prime})=1$, it holds that
@@
This definition of RDP can be easily converted to standard DP (in Equation 1) via M. While Definition 2 is unconditional on the possible set of datasets, a relaxed version of conditional RDP can be defined over a given dataset collection ${\mathcal{L}}$ such that the neighboring datasets $S,S^{\prime}\in{\mathcal{L}}$. When the dataset collection is known in advance, the conditional definition of RDP allows for tighter privacy analysis. We use this conditional version  [^1] To avoid confusion, we note that our privacy guarantees are not conditional, in the sense that it does not suffer a catastrophic failure under any dataset. in some of our analysis.
@@ 2.2 Private Mechanisms
The easiest way to inject DP guarantees in an estimation problem is to perturb the output of the non-private estimator with appropriately calibrated noise. Two most common ways to do so are Gaussian Mechanism and Laplace Mechanism. For any deterministic estimator $f$, both mechanisms add noise proportional to the global sensitivity $\Delta_{f}$, defined as the maximum difference in $f$ over all pairs of neighboring datasets. For a given privacy parameter $\varepsilon>0$, both the Gaussian mechanism, denoted ${\mathcal{M}}_{G}$, and the Laplace mechanism, denoted ${\mathcal{M}}_{L}$, produce an output of the form $f(S)+\xi$. Here, $\xi$ follows a Gaussian distribution ${\mathcal{N}}(0,\nicefrac{{\Delta_{f}^{2}}}{{\varepsilon^{2}}})$ for the Gaussian mechanism, and a Laplace distribution $Lap(\nicefrac{{\Delta_{f}}}{{\varepsilon}})$ for the Laplace mechanism.
@@ 3 Main Results
We first introduce a norm-based privacy notion, called Smooth RDP, that allows us to conduct more fine-grained analysis on the impact of pre-processing algorithms. Using this definition, we establish our main results on the privacy guarantees of a pre-processed DP pipeline.
@@ 3.1 Smooth RDP
Notation
Meaning
Mechanism
Assumptions
RDP
SRDP
$f$
Output function
${\mathcal{M}}_{G}$
$f$ is $L$-Lipschitz
$\frac{\alpha\varepsilon^{2}}{2}$
$\frac{\alpha L^{2}\tau^{2}\varepsilon^{2}}{2\Delta_{f}^{2}}$
${\mathcal{M}}_{L}$
$f$ is $L$-Lipschitz
$\varepsilon$
$\frac{L\tau}{\Delta_{f}}\varepsilon$
$Q$
Score function
${\mathcal{M}}_{E}$
$Q$ is $L$-Lipschitz
$\varepsilon$
$\frac{L\tau}{\Delta_{Q}}\varepsilon$
$\ell$
Loss function
${\mathcal{A}}_{\mathrm{GD}}$
$\ell$ is $L$-Lipschitz and $\mu$-smooth, $\sigma=\frac{L\sqrt{T}}{\varepsilon n}$
${2\alpha\varepsilon^{2}}$
$\frac{\alpha\mu^{2}\tau^{2}\varepsilon^{2}}{2L^{2}}$
$T$
Number of iteration
${\mathcal{A}}_{\mathrm{SGD-samp}}$
$\ell$ is $L$-Lipschitz and $\mu$-smooth, $\sigma=\Omega(\nicefrac{{L\sqrt{T}}}{{\varepsilon n}})$, inverse point-wise divergence $\gamma$, $1\leq\alpha\leq\min\left\{{\frac{\sqrt{T}}{\varepsilon},\frac{L^{2}T}{\varepsilon^{2}n^{2}}\log\frac{n^{2}\epsilon}{L\sqrt{T}}}\right\}$
$\frac{\alpha^{2}\varepsilon^{2}}{2}$
$\frac{\alpha\mu^{2}\tau^{2}\varepsilon^{2}\gamma^{2}}{2L^{2}}$
$\eta$
Learning rate
${\mathcal{A}}_{\mathrm{SGD-iter}}$
$\ell$ is convex, $L$-Lipschitz and $\mu$-smooth, $\sigma=\frac{8\sqrt{2\log n}\eta L}{\varepsilon\sqrt{n}}$, $\varepsilon=O(1/n\alpha^{2})$, maximum divergence $\kappa_{\tau}$, $L\sqrt{2\alpha(\alpha-1)}\leq\sigma$
$\frac{\alpha\varepsilon^{2}}{2}$
$\frac{\alpha\tau^{2}\mu^{2}n\log(n-\kappa_{\tau}+2)}{2(n-\kappa_{\tau}+1)L^{2}\log n}$
Table 1: RDP and SRDP parameters of DP mechanisms.
Our analysis on the privacy guarantees of pre-processed DP pipelines relies on a privacy notion that ensures indistinguishability between two datasets with a bounded $L_{12}$ distance. Here, the $L_{12}$ distance between two datasets $S$ and $S^{\prime}$ is defined as $d_{12}(S,S^{\prime})=\sum_{i=1}^{n}\left\lVert S_{i}-S_{i}^{\prime}\right\rVert_{2}$. We introduce this privacy notion as Smooth Rényi Differential Privacy (SRDP), defined as follows:
@@
Let $\varepsilon(\alpha,\tau)$ be a function that maps each $\alpha,\tau$ pair to a real value. A randomized algorithm ${\mathcal{A}}:{\mathcal{X}}^{n}\rightarrow{\mathbb{P}}({\mathcal{Y}})$ is $(\alpha,\varepsilon(\alpha,\tau))$-SRDP if for each $\alpha>1$ and $\tau>0$,
@@
While conditional SRDP over a set ${\mathcal{L}}$ can be considered as a special case of Pufferfish Rényi Privacy (Kifer and Machanavajjhala, 2014), we show that it satisfies desirable properties, such as sequential composition and privacy amplification by subsampling, which are not satisfied by the more general Pufferfish Renyi Privacy (Pierquin and Bellet, 2023).
@@
For DP-SGD by subsampling and iteration, the SRDP guarantee is dependent on two properties: the inverse pointwise divergence and maximum divergence of the datasets, detailed in Section A.2.2. Briefly, inverse pointwise divergence $\gamma$ measures the ratio between the $d_{12}$ distance and the maximum pointwise distance between two datasets. The maximum divergence $\kappa_{\tau}$ measures the number of points that the two datasets differ. DP-SGD with subsampling benefits from a small $\gamma$, while DP-SGD with iteration benefits from a small $\kappa_{\tau}$. As we will show in the later sections, at least one of these conditions are usually met in practice.
@@ 3.2 Privacy of Pre-Processed DP Pipelines
Before stating the main result, we introduce the term data-dependent pre-processing algorithm. A deterministic data-dependent pre-processing algorithm $\pi:{\mathcal{X}}^{n}\to{\mathcal{F}}$ takes a dataset $S$ as input and returns a pre-processing function $\pi_{S}:{\mathcal{X}}\to{\mathcal{X}}\cup\emptyset$ in a function space ${\mathcal{F}}$. Intuitively, privacy of a private algorithm is retained under non-private data-dependent pre-processing, if a single element in the dataset has bounded impact on the output of the pre-processing algorithm. We define the sensitivity of a pre-processing algorithm to quantify the impact of a simgle element below.
@@
Let $S_{1}=S\cup\{z_{1}\}$ and $S_{2}=S\cup\{z_{2}\}$ be two arbitrary neighboring datasets, we define the $L_{\infty}$ and $L_{2}$ sensitivity of a pre-processing function $\pi$ as
$\Delta_{\infty}(\pi)$ $=\sup_{\begin{subarray}{c}S_{1},S_{2}:d_{H}(S_{1},S_{2})=1\end{subarray}}d_{H}(\pi_{S_{1}}(S),\pi_{S_{2}}(S))$ (2)
$\Delta_{2}(\pi)$ $=\sup_{\begin{subarray}{c}S_{1},S_{2}:d_{H}(S_{1},S_{2})=1\end{subarray}}\max_{x\in S}\left\lVert\pi_{S_{1}}(x)-\pi_{S_{2}}(x)\right\rVert_{2}$
@@
For a set of datasets ${\mathcal{L}}$ and for any $\alpha\geq 1$, $\tau>0$, consider an algorithm ${\mathcal{A}}$ that is $(\alpha,\varepsilon(\alpha))$-RDP and $(\alpha,\tilde{\varepsilon}(\alpha,\tau))$-SRDP over the set ${\mathcal{L}}$. For a pre-processing algorithm $\pi$ with $L_{\infty}$ sensitivity $\Delta_{\infty}$ and $L_{2}$ sensitivity $\Delta_{2}$, ${\mathcal{A}}\circ\pi$ is $\left({\alpha,\widehat{\varepsilon}}\right)$-RDP over ${\mathcal{L}}$ for all $p,q\geq 1$, where
@@
While the privacy guarantee provided by Theorem 2 is conditional over a dataset collection ${\mathcal{L}}$, it can be extend to an unconditional privacy guarantee over all possible datasets using the Propose-Test-Release (PTR) framework (Dwork and Lei, 2009). We show an example of the application of PTR and its guarantees in Section 5.2.
@@
Comparison with the group privacy or DDP analysis A naive analysis on the privacy guarantee of ${\mathcal{A}}\circ\pi$ using group privacy or DDP (Zhao et al., 2017; Liu et al., 2016) provides an upper bound that grows polynomially [^2] Linearly if we consider Approximate DP. with $\Delta_{\infty}$, which can be as large as $\text{poly}(n)$. In contrast, Theorem 2 implies a tighter bound on privacy for common DP mechanisms, especially when $\Delta_{\infty}\Delta_{2}=o(n)$. [^3] Theorem 2 is also applicable for other distance metrics, as long as Smooth RDP and the sensitivity of pre-processing are defined under comparable distance metric. Next, we present various examples of pre-processing algorithms with small sensitivities $\Delta_{\infty}\Delta_{2}=O(1)$. In Figure 2, we illustrate the improvement of the privacy analysis in Theorem 2 over the conventional analysis via group privacy for numerous pre-processing algorithms.

### v4

@@ Abstract
When analyzing Differentially Private (DP) machine learning pipelines, the potential privacy cost of data-dependent pre-processing is frequently overlooked in privacy accounting. In this work, we propose a general framework to evaluate the additional privacy cost incurred by non-private data-dependent pre-processing algorithms. Our framework establishes upper bounds on the overall privacy guarantees by utilising two new technical notions: a variant of DP termed Smooth DP and the bounded sensitivity of the pre-processing algorithms. In addition to the generic framework, we provide explicit overall privacy guarantees for multiple data-dependent pre-processing algorithms, such as data imputation, quantization, deduplication, standard scaling and PCA, when used in combination with several DP algorithms. Notably, this framework is also simple to implement, allowing direct integration into existing DP pipelines.
@@ 1 Introduction
With the growing emphasis on user data privacy, Differential Privacy (DP), has become the preferred solution for safeguarding training data in Machine Learning (ML) and data analysis (Dwork et al., 2006). DP algorithms are designed to ensure that individual inputs minimally affect the algorithm’s output, thus preserving privacy. This approach is now widely adopted by various organizations for conducting analyses while maintaining user privacy (AppleDP, 2017; U.S. Census Bureau, 2020).
Pre-processing data is a standard practice in data analysis and machine learning. Techniques such as data imputation for handling missing values (Anil Jadhav & Ramanathan, 2019), deduplication for reducing memorization and eliminating bias (Kandpal et al., 2022; Lee et al., 2022), standard scaling for reducing the impact of outliers, and dimensionality reduction for denoising or visualization (Abadi et al., 2016; Zhou et al., 2021; Pinto et al., 2024) are commonly used. These pre-processings are also prevalent prior to applying DP algorithms, a pipeline that we term pre-processed DP pipeline. Among other uses, this has been shown to improve privacy-accuracy trade-off (Tramer & Boneh, 2021; Ganesh et al., 2023).
@@
There are multiple strategies to address this. A straightforward method to derive privacy guarantees for this pipeline is to use group privacy where the size of the group can be as large as the size of the dataset, thereby resulting in weak privacy guarantees. This idea, albeit not in the context of pre-processing, was previous explored under the name Dependent Differential Privacy (DDP) (Zhao et al., 2017; Liu et al., 2016). Another approach is to use public data for the pre-processing. Broadly referred to as semi-private learning algorithms, examples of these methods include pre-training on public data and learning projection functions using the public data e.g. Pinto et al. (2024); Li et al. (2022b, a); Yu et al. (2021). Despite their success, these methods crucially rely on the availability of high-quality public data.
@@
    We introduce a generic framework to quantify the privacy loss in the pre-processed DP pipeline. Applying this framework, we evaluate the impact of commonly used pre-processing techniques such as deduplication, quantization, data imputation, standard scaling, and PCA on the overall privacy guarantees.
@@
Another line of research, focusing on privacy in correlated datasets (Liu et al., 2016; Zhao et al., 2017; Humphries et al., 2023), shows that correlations in the datasets can increase privacy risk of ML models. In response, Pufferfish Differential Privacy (Kifer & Machanavajjhala, 2014; Song et al., 2017) and Dependent Differential Privacy Liu et al. (2016); Zhao et al. (2017) were proposed as privacy notions tailored for datasets with inherent dependencies. However, these definitions usually require complete knowledge of the datasets’ dependency structure or the data generating process and sometimes lead to vacuous privacy guarantees. Moreover, the application of these privacy notions usually complicates the privacy analysis, as many privacy axioms, such as composition, do not hold under these more general notions.
@@ 2.1 Rényi Differential Privacy
Differential privacy restricts the change in the output distribution of a randomized algorithm by altering a single element in the dataset. Formally, for $\varepsilon,\delta>0$, a randomized algorithm ${\mathcal{A}}$ satisfies $(\varepsilon,\delta)$-DP if for any two dataset $S,S^{\prime}$ that differ by exactly one element and any possible output of the algorithm ${\mathcal{O}}$,
${\mathbb{P}}\left[{{\mathcal{A}}(S)\in{\mathcal{O}}}\right]\leq e^{\varepsilon}{\mathbb{P}}\left[{{\mathcal{A}}(S^{\prime})\in{\mathcal{O}}}\right]+\delta.$ (1)
In this work, we mainly focus on a stronger notion of DP, known as Rényi Differential Privacy (RDP)(Mironov, 2017), based on the Rényi divergence between two distributions.
@@
Let $P,R$ be two probability distributions with $\operatorname{supp}(P)\subseteq\operatorname{supp}(R)$. Let $\alpha>1$. The Rényi divergence with order $\alpha$ between $P$ and $R$ is defined as
$D_{\alpha}(P||R)=\frac{1}{\alpha-1}\log{\mathbb{E}}_{x\sim R}\left({\frac{P(x)}{R(x)}}\right)^{\alpha}.$
Let $d_{H}(\cdot,\cdot)$ denote the Hamming distance between two datasets, we formally define RDP as follows.
@@
Let $\varepsilon(\alpha)$ be a function that maps each $\alpha$ to a positive real number. A randomized algorithm ${\mathcal{A}}$ is $(\alpha,\varepsilon(\alpha))$-RDP if for all $\alpha>1$, for any datasets $S,S^{\prime}$ differing at a single point, it holds that
@@
This definition of RDP can be easily converted to standard DP via M. While Definition 2 is unconditional on the possible set of datasets, a relaxed version of conditional RDP can be defined over a given dataset collection ${\mathcal{L}}$ such that the neighboring datasets $S,S^{\prime}\in{\mathcal{L}}$. When the dataset collection is known in advance, the conditional definition of RDP allows for tighter privacy analysis. We use this conditional version in some of our analysis [^1] To avoid confusion, we note that our privacy guarantees are not conditional, in the sense that it does not suffer a catastrophic failure under any dataset..
@@ 2.2 Private mechanisms
The easiest way to inject DP guarantees in an estimation problem is to perturb the output of the non-private estimator with appropriately calibrated noise. Two most common ways to do so are Gaussian Mechanism and Laplace Mechanism. For any deterministic estimator $f$, both mechanisms add noise proportional to the global sensitivity $\Delta_{f}$, defined as the maximum difference in $f$ over all pairs of neighboring datasets. For a given privacy parameter $\varepsilon>0$, both the Gaussian mechanism, denoted ${\mathcal{M}}_{G}$, and the Laplace mechanism, denoted ${\mathcal{M}}_{L}$, produce an output of the form $f(S)+\xi$. Here, $\xi$ follows a Gaussian distribution ${\mathcal{N}}(0,\nicefrac{{\Delta_{f}^{2}}}{{\varepsilon^{2}}})$ for the Gaussian mechanism, and a Laplace distribution $\operatorname{Lap}(\nicefrac{{\Delta_{f}}}{{\varepsilon}})$ for the Laplace mechanism.
@@ 3 Main Results
We first introduce a norm-based privacy notion, called Smooth RDP, that allows us to conduct a more fine-grained analysis on the impact of pre-processing algorithms. Using this definition, we establish our main results on the privacy guarantees of a pre-processed DP pipeline.
@@ 3.1 Smooth RDP
Table 1: RDP and SRDP parameters of DP mechanisms. We let $\gamma$ and $\kappa_{\tau}$ denote the inverse pointwise divergence and maximum divergence between two datasets.
Notation Meaning Mechanism Assumptions RDP SRDP
$f$ Output function ${\mathcal{M}}_{G}$ $f$ is $L$-Lipschitz $\frac{\alpha\varepsilon^{2}}{2}$ $\frac{\alpha L^{2}\tau^{2}\varepsilon^{2}}{2\Delta_{f}^{2}}$
${\mathcal{M}}_{L}$ $f$ is $L$-Lipschitz $\varepsilon$ $\frac{L\tau}{\Delta_{f}}\varepsilon$
$Q$ Score function ${\mathcal{M}}_{E}$ $Q$ is $L$-Lipschitz $\varepsilon$ $\frac{L\tau}{\Delta_{Q}}\varepsilon$
[…]
$\eta$ Learning rate
Our analysis on the privacy guarantees of pre-processed DP pipelines relies on a privacy notion that ensures indistinguishability between two datasets with a bounded $L_{12}$ distance. Here, the $L_{12}$ distance between two datasets $S$ and $S^{\prime}$ of size $n$ is defined as $d_{12}(S,S^{\prime})=\sum_{i=1}^{n}\left\lVert S_{i}-S_{i}^{\prime}\right\rVert_{2}$. We introduce this privacy notion as Smooth Rényi Differential Privacy (SRDP), defined as follows:
@@
Let $\varepsilon(\alpha,\tau)$ be a function that maps each $\alpha,\tau$ pair to a real value. A randomized algorithm ${\mathcal{A}}$ is $(\alpha,\varepsilon(\alpha,\tau))$-SRDP if for each $\alpha>1$ and $\tau>0$,
@@
While conditional SRDP over a set ${\mathcal{L}}$ can be considered as a special case of Pufferfish Rényi Privacy (Kifer & Machanavajjhala, 2014), we show that it satisfies desirable properties, such as sequential composition and privacy amplification by subsampling, which are not satisfied by the more general Pufferfish Renyi Privacy (Pierquin & Bellet, 2023).
@@
  • Composition Let $k$ be a positive integer. Let $\alpha>0,\varepsilon_{i}:\mathbb{R}\times\mathbb{R}\to\mathbb{R},\forall i\in[k]$. For any $i\in[k]$, if the randomized algorithms ${\mathcal{A}}_{i}$ is $(\alpha,\varepsilon_{i}(\alpha,\tau))$-SRDP over ${\mathcal{L}}$, then the composition of the $k$ algorithms $({\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k})$ is $(\alpha,\sum_{i=1}^{k}\varepsilon_{i}(\alpha,\tau))$-SRDP over ${\mathcal{L}}$.
@@
For DP-SGD by subsampling and iteration, the SRDP guarantee is dependent on two properties: the inverse pointwise divergence and maximum divergence of the datasets, detailed in Section A.2.2. Briefly, the inverse pointwise divergence $\gamma$ measures the ratio between the $d_{12}$ distance and the maximum pointwise distance between two datasets. The maximum divergence $\kappa_{\tau}$ measures the number of points that the two datasets differ. DP-SGD with subsampling benefits from a small $\gamma$, while DP-SGD with iteration benefits from a small $\kappa_{\tau}$. As we will show in the later sections, at least one of these conditions are usually met in practice.
@@ 3.2 Privacy of Pre-Processed DP Pipelines
Before stating the main result, we introduce the term data-dependent pre-processing algorithm. Let ${\mathcal{X}}\in\mathbb{R}^{d}$ be the instance space with Euclidean norm bounded by 1, i.e.$\forall x\in{\mathcal{X}}$, $\left\lVert x\right\rVert_{2}\leq 1$. A deterministic data-dependent pre-processing algorithm $\pi:{\mathcal{X}}^{n}\to{\mathcal{F}}$ takes a dataset $S$ as input and returns a pre-processing function $\pi_{S}:{\mathcal{X}}\to{\mathcal{X}}\cup\emptyset$ in a function space ${\mathcal{F}}$. Intuitively, privacy of a private algorithm is retained under non-private data-dependent pre-processing, if a single element in the dataset has bounded impact on the output of the pre-processing algorithm. We define the sensitivity of a pre-processing algorithm to quantify the impact of a simgle element below.
@@
Let $S_{1}=S\cup\{z_{1}\}$ and $S_{2}=S\cup\{z_{2}\}$ be two arbitrary neighboring datasets, we define the $L_{\infty}$ and $L_{2}$ sensitivity of a pre-processing function $\pi$ as [^2] When one of $\pi_{S_{1}}(x)$ and $\pi_{S_{2}}(x)$ is $\emptyset$, we define $\left\lVert\pi_{S_{1}}(x)-\pi_{S_{2}}(x)\right\rVert_{2}=1$.
$\Delta_{\infty}(\pi)$ $=\sup_{\begin{subarray}{c}S_{1},S_{2}:d_{H}(S_{1},S_{2})=1\end{subarray}}d_{H}(\pi_{S_{1}}(S),\pi_{S_{2}}(S)),$ (2)
$\Delta_{2}(\pi)$ $=\sup_{\begin{subarray}{c}S_{1},S_{2}:d_{H}(S_{1},S_{2})=1\end{subarray}}\max_{x\in S}\left\lVert\pi_{S_{1}}(x)-\pi_{S_{2}}(x)\right\rVert_{2}.$
@@
For a set of datasets ${\mathcal{L}}$ and for any $\alpha\geq 1$, $\tau>0$, consider an algorithm ${\mathcal{A}}$ that is $(\alpha,\varepsilon(\alpha))$-RDP and $(\alpha,\tilde{\varepsilon}(\alpha,\tau))$-SRDP over the set ${\mathcal{L}}$. For a pre-processing algorithm $\pi$ with $L_{\infty}$ sensitivity $\Delta_{\infty}$ and $L_{2}$ sensitivity $\Delta_{2}$, ${\mathcal{A}}\circ\pi$ is $\left({\alpha,\widehat{\varepsilon}}\right)$-RDP over ${\mathcal{L}}$ for all $c_{1},c_{2}\geq 1$, where
@@
$\frac{\alpha c_{2}-1}{c_{2}(\alpha-1)}\varepsilon\left({\alpha c_{2}}\right)+\tilde{\varepsilon}\left({\frac{c_{2}\alpha-1}{c_{2}-1},\Delta_{2}\Delta_{\infty}}\right)\bigg{\}}.$
@@
While the privacy guarantee provided by Theorem 2 is conditional over a dataset collection ${\mathcal{L}}$, it can be extend to an unconditional privacy guarantee over all possible datasets using the Propose-Test-Release (PTR) framework (Dwork & Lei, 2009). We show an example of the application of PTR and its guarantees in Section 5.2.
@@
Comparison with the group privacy or DDP analysis
A naive analysis on the privacy guarantee of ${\mathcal{A}}\circ\pi$ using group privacy or DDP (Zhao et al., 2017; Liu et al., 2016) provides an upper bound that grows polynomially [^3] Linearly if we consider Approximate DP. with $\Delta_{\infty}$, which can be as large as $\text{poly}(n)$. In contrast, Theorem 2 implies a tighter bound on privacy for common DP mechanisms, especially when $\Delta_{\infty}\Delta_{2}=o(n)$. [^4] Theorem 2 is also applicable for other distance metrics, as long as Smooth RDP and the sensitivity of pre-processing are defined under comparable distance metric. Next, we present various examples of pre-processing algorithms with small sensitivities $\Delta_{\infty}\Delta_{2}=O(1)$. In Figure 2, we illustrate the improvement of the privacy analysis in Theorem 2 over the conventional analysis via group privacy for numerous pre-processing algorithms.